# Resonant Multi-Scalar Production in the Generic Complex Singlet Model in the Multi-TeV Region

### v1

@@ Acknowledgements
IML would like the Pittsburgh Particle Physics Astrophysics and Cosmology center for their hospitality during the ongoing work on this paper. SDL, IML, and MS have been supported in part by the United States Department of Energy grant number DE-SC001798. SDL, MS are also supported in part by the State of Kansas EPSCoR grant program. MS is also supported in part by the United States Department of Energy under Grant Contract DE-SC0012704. SDL was also supported in part by the National Research Foundation of Korea (Grant No. NRF-2021R1A2C2009718) and by the University of Kansas General Research Funds. Digital data for the plots and benchmark points have been uploaded with the arXiv version of this paper.
@@ Appendix A Perturbative Unitarity
The perturbative unitarity matrix in Eq. 25 can be block-diagonalized by $U\mathcal{M}U^{T}$ with U given by
@@
After block-diagonalization, $\mathcal{M}$ takes the form
@@ Appendix B Bounds on Cubic and Quartic Potential Parameters
These four inequalities constrain four independent combinations of the real and imaginary parts of $e_{1}$ and $e_{2}$. The inequalities in Eqs. (66) (67) can be combined, and they become
@@
The upper bounds in Eqs. (70) and (71) are a result of placing necessary but not sufficient perturbative unitarity upper bounds on the diagonal terms of Eq. (60). Similarly, the upper bounds in Eqs. (72) and (73) can be obtained from Eq. (62). The lower bound in Eq. (70) is found by demanding the potential be bounded from below in the $\phi=S=0$ direction. Eq. (71) is from the bounded from below condition in the $\phi=A=0$ direction. Finally, the lower bounds in Eqs. (72) and (73) are found by demanding the $A=\pm S$ with $\phi=0$ directions are bounded from below.
The perturbative upper bounds here can be combined with the global minimum constraints in Eqs. (64, 65, 68, 69) to obtain a series of upper bounds on the scalar trilinear terms in the potential:
@@
$|\textrm{Re}\left(e_{1}\right)+\textrm{Re}\left(e_{2}\right)|$$\leq$ $\sqrt{96\pi}\left(m_{1}^{2}\sin^{2}\theta_{1}+m_{2}\cos^{2}\theta_{1}\right)^{1/2}$ (77)
@@
$A=0:$ $0\leq\lambda\left(d_{2}+\textrm{Re}\left(d_{1}\right)+\textrm{Re}\left(d_{3}\right)\right)-\left(\delta_{2}+\textrm{Re}\left(\delta_{3}\right)\right)^{2}$ (81)
$S=0$ $0\leq\lambda\left(d_{2}+\textrm{Re}\left(d_{1}\right)-\textrm{Re}\left(d_{3}\right)\right)-\left(\delta_{2}-\textrm{Re}\left(\delta_{3}\right)\right)^{2}$ (82)
$S=\pm A:$ $0\leq\lambda\left(d_{2}-\textrm{Re}\left(d_{1}\right)\mp\textrm{Im}\left(d_{3}\right)\right)-\left(\delta_{2}\mp\textrm{Im}\left(\delta_{3}\right)\right)^{2}$ (83)
$\phi=S=\pm A:$ $0\leq\lambda+4\left(d_{2}-\textrm{Re}\left(d_{1}\right)\mp\textrm{Im}\left(d_{3}\right)\right)+4\left(\delta_{2}\mp\textrm{Im}\left(\delta_{3}\right)\right)$ (84)
Adding together the $A=(1\pm\sqrt{2})S$ directions yields vacuum stability constraints:
ϕ=0::italic-$\phi=0:\quad\quad$ $0\leq 3\left(d_{2}+\textrm{Im}\left(d_{1}\right)\right)-2\left(\textrm{Im}\left(d_{3}\right)+\textrm{Re}\left(d_{3}\right)\right)$ (86)
ϕ0::italic-$\phi\neq 0:\quad\quad$ $0\leq 6\,\lambda\left(d_{2}+\textrm{Im}\left(d_{1}\right)\right)-4\,\lambda\left(\textrm{Im}\left(d_{3}\right)+\textrm{Re}\left(d_{3}\right)\right)$
@@
The vacuum stability bounds in Eqs. (81-83) can be combined with the perturbative unitarity bounds in Eqs. (70-73) to find:
@@
$\left|\delta_{2}\right|$$\leq$ $\sqrt{\lambda\frac{32\pi}{3}}=\frac{8}{v}\sqrt{\frac{\pi}{3}}\left(m_{1}^{2}\cos^{2}\theta_{1}+m_{2}^{2}\sin^{2}\theta_{1}\right)^{1/2}$ (89)
@@
$\begin{pmatrix}\mathcal{M}(hS\rightarrow hS)&\mathcal{M}(hS\rightarrow hA)\\ \mathcal{M}(hA\rightarrow hS)&\mathcal{M}(hA\rightarrow hA)\end{pmatrix}=\frac{1}{2}\begin{pmatrix}\delta_{2}+\textrm{Re}\left(\delta_{3}\right)&-\textrm{Im}\left(\delta\right)_{3}\\ -\textrm{Im}\left(\delta_{3}\right)&\delta_{2}-\textrm{Re}\left(\delta_{3}\right)\end{pmatrix}$ (92)
@@ Appendix C Bounds on Branching Ratios in the Multi-TeV Region
First we maximize ${\rm BR}(h_{2}\rightarrow h_{3}h_{3})$. To do this, we maximize the magnitude of $\lambda_{233}$ while minimizing the magnitudes of $\lambda_{123}$ and $\lambda_{112}$. The combinations of ${\rm Re}(e_{1}),{\rm Re}(e_{2})$ and $\delta_{2},{\rm Re}(\delta_{3})$ appear with opposite signs in $\lambda_{112}$ and $\lambda_{123}$. Hence, to get obtain larger $|\lambda_{233}|$ while keeping $|\lambda_{112}|$ smaller we choose $\textrm{Re}\left(e_{1}\right)=-\textrm{Re}\left(e_{2}\right)$ and $\delta_{2}=-\textrm{Re}\left(\delta_{3}\right)$. Since $\lambda_{233}$ is independent of ${\rm Im}(e_{1,3})$ and ${\rm Im}(\delta_{2})$, $\lambda_{123}$ can be eliminated with by $\textrm{Im}\left(e_{2}\right)=-3\textrm{Im}\left(e_{1}\right)$ and $\textrm{Im}\left(\delta_{3}\right)=0$. With these parameter choices
@@
The situation for ${\rm BR}(h_{2}\rightarrow h_{1}h_{3})$ is considerably different. First, parameters are chosen to make $\lambda_{233}$ small. In particular, while $\lambda_{112}$ is suppressed by $\sin\theta_{1}$, $\lambda_{233}$ has a term that is unsupressed by small mixing angle. Since $\lambda_{233}$ depends on different trilinear and quartic potential terms than $\lambda_{123}$, $\lambda_{233}$ can be eliminated with the choice
@@
This choice leaves $\lambda_{123}$ unchanged. Now, an upper bounds on $|\lambda_{123}|$ can be found with a combination of Eqs. (75,78,91) and
@@
To find an upper bound on the rate of of $pp\rightarrow h_{2}\rightarrow h_{1}h_{3}$, a maximum $\sin\theta_{1}$ needs to be found. Using our narrow width requirement
@@ D.1 ATLAS Higgs Signal Strengths
Since the combinations Higgs measurements in Summer 2022 ATLAS:2022vkf , ATLAS has updated their results for $h\rightarrow WW^{*}$ in the $Wh$ and $Zh$ production modes ATLAS-CONF-2022-067 from 36.1 fb-1 to 139 fb-1. Here, we describe our method of including the new results.
@@ D.2 CMS Higgs Signal Strengths
CMS has a new measurement of $h\rightarrow b\bar{b}$ in the $t\bar{t}h$ production mode CMS-PAS-HIG-19-011 from 35.9 fb-1 to 138 fb-1 since the reported Higgs combination in summer 2022 CMS:2022dwd . Similarly to the ATLAS update in the previous subsection, we provide our method of incorporating this new measurement. As with ATLAS, we first reproduce the previous CMS results. From Ref. CMS:2022dwd we have the signal strengths for different final states
@@
Using Eqs. (114,115), we find the global signal strength
@@
Using Eqs. (114,115), the new $bb$ signal strength is found to be
@@ D.3 Scalar Search Limits
Figure 9: Comparison between observed upper bounds on $\sigma(pp\rightarrow h_{2}\rightarrow h_{1}h_{1})$ reported by ATLAS from the combination ATLAS-CONF-2021-052 and individual analysis papers for (a) $4b$ ATLAS:2022hwc , (b) $bb\tau\tau$ ATLAS:2022xzm , anc (c) $bb\gamma\gamma$ ATLAS:2021ifb final state. Solid lines are for the individual channel results from the combination ATLAS-CONF-2021-052 , red dotted for the individual analysis papers ATLAS:2022hwc ; ATLAS:2022xzm ; ATLAS:2021ifb , and the blue dotted for the combination of the channels ATLAS-CONF-2021-052 .
We use the combined search for $h_{2}\rightarrow h_{1}h_{1}\rightarrow 4b,\,bb\tau\tau,\,b\gamma\gamma$ from ATLAS which appeared in the conference note ATLAS-CONF-2021-052 . There are stand alone papers for the individual searches of $4b$ ATLAS:2022hwc , $bb\tau\tau$ ATLAS:2022xzm , and $bb\gamma\gamma$ ATLAS:2021ifb with similar luminosity. Some of these searches appeared after the conference note.

### v2

@@ Acknowledgements
IML would like the Pittsburgh Particle Physics Astrophysics and Cosmology center for their hospitality during the ongoing work on this paper. SDL, IML, and MS have been supported in part by the U.S. Department of Energy Grant No. DE-SC0017988. SDL, MS are also supported in part by the State of Kansas EPSCoR grant program. MS is also supported in part by the United States Department of Energy under Grant Contract DE-SC0012704. SDL was also supported in part by the National Research Foundation of Korea (Grant No. NRF-2021R1A2C2009718) and by the University of Kansas General Research Funds. Digital data for the plots and benchmark points have been uploaded with the arXiv version of this paper.
@@ Appendix A Perturbative Unitarity
The perturbative unitarity matrix in Eq. (25) can be block diagonalized by $U\mathcal{M}U^{T}$ with $U$ given by
@@
After block diagonalization, $\mathcal{M}$ takes the form
@@ Appendix B Bounds on Cubic and Quartic Potential Parameters
These four inequalities constrain four independent combinations of the real and imaginary parts of $e_{1}$ and $e_{2}$. The inequalities in Eqs. (68,69) can be combined, and they become
@@
The upper bounds in Eqs. (72) and (73) are a result of placing necessary but not sufficient perturbative unitarity upper bounds on the diagonal terms of Eq. (62). Similarly, the upper bounds in Eqs. (74) and (75) can be obtained from Eq. (64). The lower bound in Eq. (72) is found by demanding the potential be bounded from below in the $\phi=S=0$ direction. Equation (73) is from the bounded from below condition in the $\phi=A=0$ direction. Finally, the lower bounds in Eqs. (74) and (75) are found by demanding the $A=\pm S$ with $\phi=0$ directions are bounded from below.
The perturbative upper bounds here can be combined with the global minimum constraints in Eqs. (66, 67, 70, 71) to obtain a series of upper bounds on the scalar trilinear terms in the potential:
@@
$|\textrm{Re}\left(e_{1}\right)+\textrm{Re}\left(e_{2}\right)|$$\leq$ $\sqrt{96\pi}\left(m_{1}^{2}\sin^{2}\theta_{1}+m^{2}_{2}\cos^{2}\theta_{1}\right)^{1/2}$ (79)
@@
$z=0:$ $0\leq\lambda\left(d_{2}+\textrm{Re}\left(d_{1}\right)+\textrm{Re}\left(d_{3}\right)\right)-\left(\delta_{2}+\textrm{Re}\left(\delta_{3}\right)\right)^{2}$ (83)
$y=0:$ $0\leq\lambda\left(d_{2}+\textrm{Re}\left(d_{1}\right)-\textrm{Re}\left(d_{3}\right)\right)-\left(\delta_{2}-\textrm{Re}\left(\delta_{3}\right)\right)^{2}$ (84)
$y=\pm z:$ $0\leq\lambda\left(d_{2}-\textrm{Re}\left(d_{1}\right)\mp\textrm{Im}\left(d_{3}\right)\right)-\left(\delta_{2}\mp\textrm{Im}\left(\delta_{3}\right)\right)^{2}$ (85)
$x=y=\pm z:$ $0\leq\lambda+4\left(d_{2}-\textrm{Re}\left(d_{1}\right)\mp\textrm{Im}\left(d_{3}\right)\right)+4\left(\delta_{2}\mp\textrm{Im}\left(\delta_{3}\right)\right)$ (86)
Adding together the $z=(1\pm\sqrt{2})y$ directions yields vacuum stability constraints:
$x=0:\quad\quad$ $0\leq 3\left(d_{2}+\textrm{Im}\left(d_{1}\right)\right)-2\left(\textrm{Im}\left(d_{3}\right)+\textrm{Re}\left(d_{3}\right)\right)$ (88)
$x\neq 0:\quad\quad$ $0\leq 6\,\lambda\left(d_{2}+\textrm{Im}\left(d_{1}\right)\right)-4\,\lambda\left(\textrm{Im}\left(d_{3}\right)+\textrm{Re}\left(d_{3}\right)\right)$
@@
The vacuum stability bounds in Eqs. (83-85) can be combined with the perturbative unitarity bounds in Eqs. (72-75) to find:
@@
$\left|\delta_{2}\right|$$\leq$ $\sqrt{\lambda\frac{32\pi}{3}}=\frac{8}{v_{\rm EW}}\sqrt{\frac{\pi}{3}}\left(m_{1}^{2}\cos^{2}\theta_{1}+m_{2}^{2}\sin^{2}\theta_{1}\right)^{1/2}$ (91)
@@
$\begin{pmatrix}\mathcal{M}(hS\rightarrow hS)&\mathcal{M}(hS\rightarrow hA)\\ \mathcal{M}(hA\rightarrow hS)&\mathcal{M}(hA\rightarrow hA)\end{pmatrix}=\frac{1}{2}\begin{pmatrix}\delta_{2}+\textrm{Re}\left(\delta_{3}\right)&-\textrm{Im}\left(\delta_{3}\right)\\ -\textrm{Im}\left(\delta_{3}\right)&\delta_{2}-\textrm{Re}\left(\delta_{3}\right)\end{pmatrix}$ (94)
@@ Appendix C Bounds on Branching Ratios in the Multi-TeV Region
First we maximize ${\rm BR}(h_{2}\rightarrow h_{3}h_{3})$. To do this, we maximize the magnitude of $\lambda_{233}$ while minimizing the magnitudes of $\lambda_{123}$ and $\lambda_{112}$. The combinations of ${\rm Re}(e_{1}),{\rm Re}(e_{2})$ and $\delta_{2},{\rm Re}(\delta_{3})$ appear with opposite signs in $\lambda_{112}$ and $\lambda_{123}$. Hence, to obtain larger $|\lambda_{233}|$ while keeping $|\lambda_{112}|$ smaller we choose $\textrm{Re}\left(e_{1}\right)=-\textrm{Re}\left(e_{2}\right)$ and $\delta_{2}=-\textrm{Re}\left(\delta_{3}\right)$. Since $\lambda_{233}$ is independent of ${\rm Im}(e_{1,3})$ and ${\rm Im}(\delta_{2})$, $\lambda_{123}$ can be eliminated with by $\textrm{Im}\left(e_{2}\right)=-3\textrm{Im}\left(e_{1}\right)$ and $\textrm{Im}\left(\delta_{3}\right)=0$. With these parameter choices
@@
The situation for ${\rm BR}(h_{2}\rightarrow h_{1}h_{3})$ is considerably different. First, parameters are chosen to make $\lambda_{233}$ small. In particular, while $\lambda_{112}$ is suppressed by $\sin\theta_{1}$, $\lambda_{233}$ has a term that is unsuppressed by small mixing angle. Since $\lambda_{233}$ depends on different trilinear and quartic potential terms than $\lambda_{123}$, $\lambda_{233}$ can be eliminated with the choice
@@
This choice leaves $\lambda_{123}$ unchanged. Now, an upper bound on $|\lambda_{123}|$ can be found with a combination of Eqs. (77,80,93) and
@@
To find an upper bound on the rate of $pp\rightarrow h_{2}\rightarrow h_{1}h_{3}$, a maximum $\sin\theta_{1}$ needs to be found. Using our narrow width requirement
@@ D.1 ATLAS Higgs Signal Strengths
Since the combinations of Higgs measurements in summer 2022 ATLAS:2022vkf , ATLAS has updated their results for $h\rightarrow WW^{*}$ in the $Wh$ and $Zh$ production modes ATLAS-CONF-2022-067 from 36.1 fb-1 to 139 fb-1. Here, we describe our method of including the new results.
@@ D.2 CMS Higgs Signal Strengths
CMS has a new measurement of $h\rightarrow b\bar{b}$ in the $t\bar{t}h$ production mode CMS-PAS-HIG-19-011 from 35.9 fb-1 to 138 fb-1 since the reported Higgs combination in summer 2022 CMS:2022dwd . Similar to the ATLAS update in the previous subsection, we provide our method of incorporating this new measurement. As with ATLAS, we first reproduce the previous CMS results. From Ref. CMS:2022dwd we have the signal strengths for different final states
@@
Using Eqs. (116,117), we find the global signal strength
@@
Using Eqs. (116,117), the new $bb$ signal strength is found to be
@@ D.3 Scalar Search Limits
Figure 9: Comparison between observed upper bounds on $\sigma(pp\rightarrow h_{2}\rightarrow h_{1}h_{1})$ reported by ATLAS from the combination ATLAS-CONF-2021-052 and individual analysis papers for (a) $4b$ ATLAS:2022hwc , (b) $bb\tau\tau$ ATLAS:2022xzm , and (c) $bb\gamma\gamma$ ATLAS:2021ifb final state. Solid lines are for the individual channel results from the combination ATLAS-CONF-2021-052 , red dotted for the individual analysis papers ATLAS:2022hwc ; ATLAS:2022xzm ; ATLAS:2021ifb , and the blue dotted for the combination of the channels ATLAS-CONF-2021-052 .
We use the combined search for $h_{2}\rightarrow h_{1}h_{1}\rightarrow 4b,\,bb\tau\tau,\,b\gamma\gamma$ from ATLAS which appeared in the conference note ATLAS-CONF-2021-052 . There are stand-alone papers for the individual searches of $4b$ ATLAS:2022hwc , $bb\tau\tau$ ATLAS:2022xzm , and $bb\gamma\gamma$ ATLAS:2021ifb with similar luminosity. Some of these searches appeared after the conference note.